A trader creates a long butterfly spread from options with strike prices $60, $65, and $70 by trading a total of 400 options. The options are worth $11, $14, and $18. What is the maximum net gain (after the cost of the options is taken into account)? Show all working in detail including the strategy, cost and payoff of the options and total payoff.
1
Expert's answer
2015-11-03T00:00:47-0500
When the strike price is $60: the net gain is NG = (60 - 11)*400 = $19,600
When the strike price is $65: the net gain is NG = (65 - 14)*400 = $20,400
When the strike price is $70: the net gain is NG = (70 - 18)*400 = $20,800
So, the highest net gain is $20,800 when the strike price is 70$.